M 408C 58985,58990,58995 Fall ’06 TTh 12:30-2:00 ECJ 1.202; MW 8-9 RLM 6.124 (58985), MW 3-4 ESB 133 (58990), MW 4-5 RLM 5.122 (58995)

*** T Th classes will be lectures, M W discussion sections.

Instructor: C. Friedman RLM 11.140 471-5161 office hrs: T Th 2:30-3:30 or by appt. Web Page (homework assignments are listed): www.ma.utexas.edu/˜ friedman

Text and Syllabus: Calculus, James Stewart 5th ed. Syllabus; 2.1-2.6, 3.1-3.10, 4.1-4.7, 4.10, 5.1-5.5, 6.1-6.2, 6.5, 7.1, 7.2*-7.4*, 7.5, 8.1-8.5 (There may be some changes in the syllabus.)

Exams: We will have exams on 10/5 (class 11; Thurs.), 11/9 (class 21; Thurs.), 12/7 (Class 28; Thurs.) (No “Final Exam” planned [subject to Department approval])

* Exam grades will be scaled (if the unscaled average is ≤ 75) to some value in the 75 to 80 range. MAKEUPS WILL BE GIVEN VERY RELUCTANTLY

Homework: Homework will be assigned each Tuesday in class and will be due the next Monday in Discussion Section. A serious effort should be made on these exercises. It will probably be difficult to completely learn the course material without working problems; also exam questions may bear some resemblance to homework exercises. Additionally, all homework together will be counted as a grade equivalent to one exam grade.

Quizes: Three quizes will be given in the discussion classes on 9/25 (Mon.), 10/30 (Mon.), and 11/29 (Wed.). These will probably be short (about 1/2 hour) and consist of one problem. All three quiz grades will be combined to form one grade counted equal to a single test grade.

Final Grades: In computing a final grade, I will count the three tests, one total homework grade and one total quiz grade equally. (So there will be five grades to average.)